The perimeter of a standard-sized rectangular rug is 28 ft. The length is 2 ft. longer than the width. find the dimensions
Answer by Flannery(124)

Perimiter is two times width plus two times length.
Width is x.
Length is x+2
28=2x+2(x+2)
28=2x+2x+4
28=4x+4
24=4x
24/4=6
so x=6
28=2(6)+2(6+2)
28=12+12+2
28=28
Solution checks.